[Brahms-l] Shift report 20050428 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Thu Apr 28 2005 - 00:16:13 EDT
Shifter leader was FR
Shifters were 


 - No beam during this shift. Problems with Cryo lead flow, Booster
   magnet power supply and power dips, and then with the weather!
 - Only pedestal runs: 14312,14316
 - Problem with D1 magnet, could'nt turn it on, called the expert.


_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 28 00:16:30 2005

This archive was generated by hypermail 2.1.8 : Thu Apr 28 2005 - 00:16:47 EDT